Charlotte's Crochet Shoppe has 16,100 shares of common stock outstanding at a price per share of $81 and a rate of return of 11.85 percent. The company also has 340 bonds outstanding, with a par value of $2,000 per bond. The pretax cost of debt is 6.25 percent and the bonds sell for 99 percent of par. What is the firm's WACC if the tax rate is 35 percent?
Answer #1

Market value of shares = 16100*81 = 1304100

Market value of bonds = 340*2000*99% = 673200

Total value = 1977300

Cost of equity = 11.85%

Cost of bonds: 6.25%*(1-0.35) = 4.0625%

WACC = 11.85%*1304100/1977300 + 4.0625%*673200/1977300

=9.2%